a. III- (5pts) In an econometrics course, the relationship between the expected score earned by students s related to the timplease answer this economics question completely..thank you its urgent.
Answer #1

##According to Chegg policy it is not possible to answer multiple questions -Questions with sub questions are fine.

a) Change in estimated score = change in time* slope

Change in estimated score= (10)0.24 = 2.4

So when time given increases by 10 minutes ,the score would increase by 2.4 points

b) We have the regression function

E(score) = 49+0.24*time

Putting time = 120 minutes

E(score) = 49+0.24(120)= 77.8 points

c) Actual score is 90

Estimated score is E(score)= 49+0.24*time

Residual = Actual - predicted

Residual= 90-49-0.24*time = 41-0.24*time

d) E(score )= 49+0.24*(140)

E(average score) = 82.6

e) In part d we are given the average score and time , the score varies from student to student , but the average is 140